# A Coherence-Based Measure of AGI

### v1

@@ Abstract
Recent work by Hendrycks et al. (2025) formalized Artificial General Intelligence (AGI) as the arithmetic mean of proficiencies across cognitive domains derived from the Cattell–Horn–Carroll (CHC) model of human cognition. While elegant, this definition assumes compensability—that exceptional ability in some domains can offset failure in others. True general intelligence, however, should reflect coherent sufficiency: balanced competence across all essential domains. We propose a coherence-aware measure of AGI based on the integral of generalized means over a continuum of compensability exponents. This formulation spans arithmetic, geometric, and harmonic regimes, and the resulting area under the curve (AUC) quantifies robustness under varying compensability assumptions. Unlike the arithmetic mean, which rewards specialization, the AUC penalizes imbalance and captures inter-domain dependency. Applied to published CHC-based domain scores for GPT-4 and GPT-5, the coherence-adjusted AUC reveals that both systems remain far from general competence despite high arithmetic scores (e.g., GPT-5 at 24%). Integrating the generalized mean thus yields a principled, interpretable, and stricter foundation for measuring genuine progress toward AGI.
@@ 1 Introduction
Artificial General Intelligence (AGI) is often described as the ultimate goal of artificial intelligence research—the creation of systems with intelligence comparable to that of humans. Yet, despite decades of progress, measuring advancement toward this goal remains elusive, largely because both intelligence and generality lack rigorous operational definitions (Chollet, 2019).
@@
Recent work by Hendrycks et al. (2025) takes a decisive step toward formalizing AGI. Building on the informal notion that “AGI is an AI that can match or exceed the cognitive versatility and proficiency of a well-educated adult,” their framework grounds evaluation in human psychometrics through the Cattell–Horn–Carroll (CHC) theory of cognitive abilities (Carroll, 1993; McGrew, 2009, 2023a)—the dominant empirical model of human intelligence. By decomposing cognition into ten broad domains (e.g., reasoning, memory, perception, and speed) and averaging performance across them, they define a unified AGI score as the arithmetic mean of domain proficiencies.
@@
While elegant, this formulation implicitly assumes a degree of compensability: that exceptional performance in some faculties can offset severe deficiencies in others. As a result, an AI system could appear “general” while failing entirely in critical domains such as reasoning or memory. This assumption conflicts with both psychometric evidence and systems theory. In human cognition, abilities are interdependent—reasoning depends on working memory, perception constrains abstraction, and learning relies on durable long-term memory (Carroll, 1993; McGrew, 2009, 2023a). Empirical studies in cognitive psychology consistently find that extreme imbalance across faculties is associated with functional impairment rather than high general intelligence (Cattell, 1987). Likewise, in complex engineered systems, overall capability is limited by the weakest component—a “bottleneck” or limiting-factor dynamic (Keeney and Raiffa, 1993; Kitano, 2004), also acknowledged by Hendrycks et al. (2025).
@@
We argue that general intelligence should instead reflect coherent sufficiency: balanced competence across all essential faculties without catastrophic failure in any. This principle resonates across disciplines—psychometrics emphasizes coherence among subtests Carroll (1993); McGrew (2009), systems theory highlights limiting-factor dynamics Kitano (2004), and multi-criteria decision analysis formalizes non-compensatory aggregation Keeney and Raiffa (1993). Collectively, these perspectives converge on the view that true generality arises from robustness and systemic balance rather than isolated excellence.
@@
To formalize this intuition, we generalize the arithmetic aggregation proposed by Hendrycks et al. (2025) into the continuous family of generalized means (Bullen, 2013) parameterized by a compensability exponent $p$. When $p=1$, the formulation reduces to the arithmetic mean; as $p$ decreases, it increasingly penalizes imbalance among cognitive domains, transitioning toward non-compensatory evaluation. Integrating the generalized mean over a range of $p$ values yields an area-under-the-curve (AUC) score that captures the stability of performance under varying compensability assumptions. This formulation preserves continuity with prior definitions while introducing a coherence-sensitive and theoretically interpretable measure of generality.
@@
    Conceptual: We identify compensability as a key limitation in current CHC-based AGI definitions and connect it to broader theories of cognitive coherence and inter-domain dependency.
@@
    Methodological: We propose measuring AGI as the integral of generalized means across compensability exponents, producing a single AUC-based indicator of systemic coherence.
@@
    Empirical: We apply this framework to published AGI domain scores for GPT-4 and GPT-5, showing that both models’ apparent generality collapses under coherence-based evaluation, revealing persistent structural imbalances.
By reframing AGI evaluation around coherence rather than compensation, this work offers a stricter, more interpretable foundation for tracking genuine progress toward general intelligence.
@@ 2.1 Psychometric Models of General Intelligence
Human intelligence is widely understood as a structured hierarchy of interdependent abilities rather than a single undifferentiated trait. The Cattell–Horn–Carroll (CHC) theory of cognitive abilities (Carroll, 1993; McGrew, 2009, 2023a) provides the dominant empirical framework for modeling this structure. CHC organizes cognition into a three-stratum hierarchy: a broad general intelligence factor ($g$) at the apex, a set of broad abilities (e.g., fluid reasoning, crystallized knowledge, memory, and processing speed) beneath it, and dozens of narrow abilities at the base. Over the past three decades, nearly all major human IQ tests have adopted the CHC approach for test construction and interpretation (Keith and Reynolds, 2010; McGrew, 2023b).
@@ 2.2 CHC-Inspired AGI Evaluation
Most AI benchmarks to date have assessed narrow capabilities in isolation, such as ImageNet for visual recognition (Deng et al., 2009), MATH for mathematical reasoning (Hendrycks et al., 2021), PIQA for commonsense inference (Bisk et al., 2020). Such benchmarks measure specialization, not generality. To address this, Hendrycks et al. (2025) proposed the first psychometrically grounded definition of AGI, explicitly mapping CHC’s cognitive domains to artificial analogues. Their framework decomposes machine cognition into ten broad abilities—General Knowledge, Reading and Writing, Mathematics, On-the-Spot Reasoning, Working Memory, Long-Term Memory (Storage and Retrieval), Visual Processing, Auditory Processing, and Speed—and evaluates proficiency in each through adapted psychometric tasks.
These domain scores are aggregated using the arithmetic mean, producing a single AGI Score between 0% and 100%. For instance, Hendrycks et al. (2025) report scores of 27% for GPT-4 (Achiam et al., 2023) and 58% for GPT-5, illustrating both rapid improvement and persistent gaps relative to human-level cognition. However, this additive aggregation assumes, at least implicitly, that cognitive abilities are independent and somehow compensatory, i.e, that exceptional performance in one domain can hide complete absence in another. As a result, the arithmetic mean can overstate overall competence, producing deceptively high composite scores for systems that appear broadly capable while still exhibiting catastrophic deficits in essential faculties such as long-term memory storage or cross-modal reasoning.
Table 1: Domain-level AGI scores reported by Hendrycks et al. (2025). K = General Knowledge; RW = Reading and Writing; M = Mathematics; R = On-the-Spot Reasoning; WM = Working Memory; MS = Long-Term Memory Storage; MR = Long-Term Memory Retrieval; V = Visual Processing; A = Auditory Processing; S = Speed.
@@ 2.3 Compensability in Complex Systems
In multi-criteria decision theory (Keeney and Raiffa, 1993), compensatory aggregation methods assume that trade-offs between dimensions are permissible—strong performance in one criterion can offset weakness in another. By contrast, non-compensatory aggregation models treat dimensions as jointly constraining, such that poor performance on any critical dimension limits the overall outcome. This framing parallels the limiting-factor principle and analyses of robustness in complex engineered systems (Kitano, 2004), both of which emphasize that system-level functioning can be bounded by its weakest component.
In psychometrics, interdependence among cognitive abilities has been repeatedly observed. Factor-analytic studies within the CHC framework reveal a pervasive positive manifold—individuals who perform well in one broad ability tend to perform well across others (Carroll, 1993; McGrew, 2009, 2023a). This pattern implies mutual reinforcement among abilities rather than independence. Furthermore, empirical work shows that limitations in core processes such as working memory or processing speed can constrain higher-level reasoning and learning performance (Fry and Hale, 2000). Accordingly, simple averages across domains may obscure such dependencies, overstating general competence when specific foundational faculties are weak.
2.4 Motivation for a Coherent AGI Metric
Given this background, an AGI metric should not merely sum capabilities but ensure their coherent sufficiency—a minimal level of competence across all essential faculties. An AI that fails completely in long-term memory or perception cannot be considered “general,” regardless of its average score. The arithmetic mean used by Hendrycks et al. (2025) fails to capture this constraint. In contrast, non-compensatory aggregation naturally enforces balance by penalizing uneven or “brittle” cognitive profiles. The generalized mean offers a principled mathematical family that interpolates between compensatory and non-compensatory regimes through a single parameter $p$. This property motivates our proposed measure in the following section, which redefines general intelligence as balanced, interdependent competence across all CHC-inspired cognitive domains.
3 A Coherence-Based Measure of AGI
Let $s_{i}\in[0,1]$ denote the normalized proficiency of an AI system in cognitive domain $i$, for $i=1,\dots,n$. Hendrycks et al. (2025) operationalize overall general intelligence as the arithmetic mean of these domain scores:
@@ 3 A Coherence-Based Measure of AGI
This formulation, denoted as $\mathrm{AGI}_{1}$, assumes that strengths in some faculties can compensate for weaknesses in others—a property known as compensability. While convenient for summarizing benchmark performance, such compensatory aggregation is problematic for evaluating general intelligence: substantial deficits in even a single faculty (e.g., reasoning, memory, or perception) can impose systemic limitations that invalidate apparent gains elsewhere. Consequently, $\mathrm{AGI}_{1}$ may overestimate holistic capability by averaging over structurally dependent abilities that, in practice, interact and constrain one another.
Generalized Mean.
@@
To relax the fully compensatory assumption of the arithmetic mean, we extend the arithmetic mean to the standard power-mean (or generalized-mean) family parameterized by an exponent $p$ (Bullen, 2013). Each value of $p\in\mathbb{R}$ corresponds to a distinct coherence regime for general intelligence, controlling the degree to which strengths can offset weaknesses:
@@
$\mathrm{AGI}_{p}=\begin{cases}\biggl(\dfrac{1}{n}\sum_{i=1}^{n}\max(s_{i},\varepsilon)^{p}\biggr)^{\!1/p},&p\neq 0,\\[8.0pt] \biggl(\prod_{i=1}^{n}\max(s_{i},\varepsilon)\biggr)^{\!1/n},&p=0~\text{(geometric mean).}\end{cases}$ (2)
Here, $\varepsilon>0$ is a small floor constant introduced to prevent numerical collapse when any $s_{i}=0$. It can be interpreted as a minimal competence threshold or a noise floor in domain measurement. In practice, $\varepsilon$ defines the lowest detectable ability and ensures stable aggregation in the non-compensatory regime. In all experiments, $\varepsilon$ was fixed at $10^{-6}$, a value sufficiently small to avoid affecting normalized proficiencies while guaranteeing numerical stability.
Interpretation of $p$.
The exponent $p$ regulates the degree of compensability among domains:
  • $p=1$: arithmetic mean—strongly compensatory regime in Hendrycks et al. (2025);
  • $p=0$: geometric mean—moderate coupling, penalizing imbalance;
  • p=1p=-1: harmonic mean—strongly non-compensatory, emphasizing the weakest abilities;
  • $p\to$-$\infty$: minimum score—a strict bottleneck or “weakest-link” formulation.
@@
As $p$ decreases, the evaluation transitions from average proficiency toward coherent sufficiency, reflecting the intuition that general intelligence requires balanced competence across all essential domains rather than isolated strengths.
Taking the definition of $\mathrm{AGI}_{p}$ to the limit as $p\to$-$\infty$ yields a fully non-compensatory formulation that effectively reduces to the minimum domain score. This represents the strictest notion of coherence among all $\mathrm{AGI}_{p}$ measures. However, such an extreme is impractical as an indicator of progress: it is overly sparse and constant to improvements when the weakest domain remains unchanged. A useful evaluation signal should therefore remain sensitive to gradual gains—while still penalizing imbalance. Even the harmonic mean (p=1p=-1) exhibits high sparsity (Table 2), whereas $p=-\tfrac{1}{2}$ already imposes substantial penalization. Conversely, the arithmetic mean ($p=1$) is overly permissive, often producing inflated aggregate scores (e.g., $50\%$) even when several domains fall below $50\%$ proficiency and at least one approaches zero (see GPT-5 in Table 1). These observations motivate restricting $p$ to the moderate interval [1,1][-1,1]: values below $1$ enforce non-compensatory behavior, while avoiding excessively small $p$ prevents signal collapse. The geometric mean ($p=0$) thus serves as a balanced “sweet spot” between sensitivity and non-compensability, and integrating $\mathrm{AGI}_{p}$ across a range of $p$ values further refines this principle—capturing robustness, sensitivity, and systemic coherence within a single scalar measure.
Aggregate robustness.

### v2

@@ Abstract
Recent approaches to evaluating Artificial General Intelligence (AGI) typically summarize a system’s capability using the arithmetic mean of its proficiencies across multiple cognitive domains. While simple, this implicitly assumes compensability: exceptional performance in some areas can offset severe deficiencies in others. Genuine general intelligence, however, requires coherent sufficiency: balanced competence across all essential faculties. We introduce a coherence-based measure of AGI that integrates the generalized mean over a continuum of compensability exponents. This yields an area-under-the-curve (AUC) metric spanning arithmetic, geometric, and harmonic regimes, quantifying how robust an evaluated capability remains as compensability assumptions become stricter. Unlike the arithmetic mean, which rewards specialization, the AUC penalizes imbalance and exposes bottlenecks that constrain performance. To illustrate the framework, we apply it to cognitive profiles derived from the Cattell–Horn–Carroll (CHC) model, showing how coherence-based aggregation highlights imbalances that are obscured by arithmetic averaging. As a second, independent example, we apply the same methodology to a set of 17 heterogeneous benchmarks, demonstrating how coherence-based evaluation can reveal unevenness even in narrower task collections. These examples show that the proposed approach offers a principled, interpretable, and stricter foundation for measuring progress toward AGI.
@@ 1 Introduction
Artificial General Intelligence (AGI) is often described as the ultimate goal of Artificial Intelligence (AI) research, the creation of systems with intelligence comparable to that of humans. Yet, despite decades of progress, measuring advancement toward this goal remains elusive, largely because both intelligence and generality lack rigorous operational definitions (chollet2019measure). While individual capabilities can be evaluated, their integration into a single indicator of AGI is far less straightforward.
@@
Recent work by hendrycks2025agidefinition builds on the informal notion that “AGI is an AI that can match or exceed the cognitive versatility and proficiency of a well-educated adult,” their framework grounds evaluation in human psychometrics through the Cattell–Horn–Carroll (CHC) theory of cognitive abilities (carroll1993human; mcgrew2009chc; mcgrew2023carroll), the dominant empirical model of human intelligence. By decomposing cognition into ten broad domains (e.g., reasoning, memory, perception, and speed) and averaging performance across them, they define a unified AGI score as the arithmetic mean of domain proficiencies.
@@
While simple and appealing, this formulation, for example, implicitly assumes a degree of compensability: that exceptional performance in some faculties can offset severe deficiencies in others. As a result, an AI system could appear “general” while failing entirely in critical domains such as reasoning or memory. This assumption conflicts with both psychometric evidence and systems theory. In human cognition, abilities are interdependent, reasoning depends on working memory, perception constrains abstraction, and learning relies on durable long-term memory (carroll1993human; mcgrew2009chc; mcgrew2023carroll). Empirical studies in cognitive psychology consistently find that extreme imbalance across faculties is associated with functional impairment rather than high general intelligence (cattell1987intelligence). Likewise, in complex engineered systems, overall capability is limited by the weakest component, a “bottleneck” or limiting-factor dynamic (keeney1993decisions; kitano2004biological), also acknowledged by hendrycks2025agidefinition.
@@
We argue that general intelligence should instead reflect coherent sufficiency: consistently high competence across all essential faculties, such that no domain falls to a low level that would preclude genuine generality. This principle resonates across disciplines, psychometrics emphasizes coherence among subtests carroll1993human; mcgrew2009chc, systems theory highlights limiting-factor dynamics kitano2004biological, and multi-criteria decision analysis formalizes non-compensatory aggregation keeney1993decisions. Collectively, these perspectives converge on the view that true generality arises from robustness and systemic balance rather than isolated excellence.
@@
To formalize this intuition, we extend the arithmetic aggregation to the continuous family of generalized means (bullen2013handbook), parameterized by a compensability exponent $p$. When $p=1$, the measure reduces to the arithmetic mean; as $p$ decreases, it increasingly penalizes imbalance across cognitive domains, smoothly transitioning toward non-compensatory evaluation. The resulting curve over $p$ (e.g., Figure 2) provides a concise diagnostic of coherence: models with flatter, higher curves exhibit more uniformly distributed competence and greater robustness to compensability assumptions.
Integrating this curve over a specified range of $p$ yields an area-under-the-curve (AUC) score, expressed as a percentage, that summarizes the stability of a system’s performance under increasingly strict compensability assumptions. This construction preserves continuity with prior arithmetic-mean definitions while introducing a coherence-based and theoretically grounded measure of generality.
@@
    Conceptual: We identify compensability as a fundamental limitation of existing AGI measurement practices and situate it within broader theories of cognitive coherence, systems interdependence, and limiting-factor dynamics.
@@
    Methodological: We propose evaluating AGI through the full AGIp curve, characterizing how performance behaves across compensability regimes, and through its integral, an AUC-based indicator that captures the coherence and robustness of a system’s capabilities.
@@
    Illustrative: We validate the framework in two independent settings: (i) CHC-style domain profiles for GPT-4 and GPT-5, and (ii) a heterogeneous suite of 17 benchmark scores for Gemini 3 Pro, GPT-5.1, Gemini 2.5 Pro, and Claude Sonnet 4.5. In both analyses, coherence-based aggregation exposes capability imbalances and limiting factors that arithmetic means obscure, demonstrating the practical value of the approach.
By reframing AGI evaluation around coherence rather than compensation, this work provides a stricter, more interpretable, and more principled criterion for assessing genuine progress toward general intelligence.
Figure 1: Comparison of model performance across aggregation exponents pp. Curves show $\mathrm{AGI}_{p}$ values; shaded regions indicate the AUC.
@@ 2.1 Psychometric Models of General Intelligence
Human intelligence is widely understood as a structured hierarchy of interdependent abilities rather than a single undifferentiated trait. The Cattell–Horn–Carroll (CHC) theory of cognitive abilities (carroll1993human; mcgrew2009chc; mcgrew2023carroll) provides the dominant empirical framework for modeling this structure. CHC organizes cognition into a three-stratum hierarchy: a broad general intelligence factor ($g$) at the apex, a set of broad abilities (e.g., fluid reasoning, crystallized knowledge, memory, and processing speed) beneath it, and dozens of narrow abilities at the base. Over the past three decades, nearly all major human IQ tests have adopted the CHC approach for test construction and interpretation (keith2010cattell; mcgrew2023chc).
@@ 2.2 CHC-Inspired AGI Evaluation
Most AI benchmarks to date have assessed narrow capabilities in isolation, such as ImageNet for visual recognition (deng2009imagenet), MATH for mathematical reasoning (hendrycks2021measuring), PIQA for commonsense inference (bisk2020piqa). Such benchmarks measure specialization, not generality. To address this, hendrycks2025agidefinition proposed the first psychometrically grounded definition of AGI, explicitly mapping CHC’s cognitive domains to artificial analogues. Their framework decomposes machine cognition into ten broad abilities—General Knowledge, Reading and Writing, Mathematics, On-the-Spot Reasoning, Working Memory, Long-Term Memory (Storage and Retrieval), Visual Processing, Auditory Processing, and Speed—and evaluates proficiency in each through adapted psychometric tasks.
These domain scores are aggregated using the arithmetic mean, producing a single AGI Score between 0% and 100%. For instance, hendrycks2025agidefinition report scores of 27% for GPT-4 (achiam2023gpt) and 58% for GPT-5, illustrating both rapid improvement and persistent gaps relative to human-level cognition. However, this additive aggregation assumes, at least implicitly, that cognitive abilities are independent and somehow compensatory, i.e, that exceptional performance in one domain can hide complete absence in another. As a result, the arithmetic mean can overstate overall competence, producing deceptively high composite scores for systems that appear broadly capable while still exhibiting catastrophic deficits in essential faculties such as long-term memory storage or cross-modal reasoning.
Table 1: Domain-level AGI scores reported by hendrycks2025agidefinition. K = General Knowledge; RW = Reading and Writing; M = Mathematics; R = On-the-Spot Reasoning; WM = Working Memory; MS = Long-Term Memory Storage; MR = Long-Term Memory Retrieval; V = Visual Processing; A = Auditory Processing; S = Speed.
@@ 2.3 Compensability in Complex Systems
In multi-criteria decision theory (keeney1993decisions), compensatory aggregation methods assume that trade-offs between dimensions are permissible—strong performance in one criterion can offset weakness in another. By contrast, non-compensatory aggregation models treat dimensions as jointly constraining, such that poor performance on any critical dimension limits the overall outcome. This framing parallels the limiting-factor principle and analyses of robustness in complex engineered systems (kitano2004biological), both of which emphasize that system-level functioning can be bounded by its weakest component.
In psychometrics, interdependence among cognitive abilities has been repeatedly observed. Factor-analytic studies within the CHC framework reveal a pervasive positive manifold—individuals who perform well in one broad ability tend to perform well across others (carroll1993human; mcgrew2009chc; mcgrew2023carroll). This pattern implies mutual reinforcement among abilities rather than independence. Furthermore, empirical work shows that limitations in core processes such as working memory or processing speed can constrain higher-level reasoning and learning performance (fry2000relationships). Accordingly, simple averages across domains may obscure such dependencies, overstating general competence when specific foundational faculties are weak.
2.4 Motivation for a Coherent-Based AGI Metric
Given this background, an AGI metric should not merely sum capabilities but ensure their coherent sufficiency, a minimal level of competence across all essential faculties. An AI that fails completely in long-term memory or perception cannot be considered “general,” regardless of its average score. In contrast, non-compensatory aggregation naturally enforces balance by penalizing uneven or “brittle” cognitive profiles. The generalized mean offers a principled mathematical family that interpolates between compensatory and non-compensatory regimes through a single parameter $p$. This property motivates our proposed measure in the following section, which redefines general intelligence as balanced, interdependent competence across all CHC-inspired cognitive domains.
3 A Coherence-Based Measure of General Intelligence
Evaluating progress toward AGI requires aggregating performance across multiple capabilities that jointly determine a system’s generality. Let
$s_{i}\in[0,100],\qquad i=1,\dots,n,$
denote the normalized proficiencies of a model across $n$ benchmark dimensions. These may represent cognitive domains, task clusters, or heterogeneous evaluations drawn from diverse modalities. Our framework makes no assumptions about the underlying ontology: each $s_{i}$ need only reflect a meaningful ability that contributes to general-purpose competence.
A widely used aggregation rule, adopted for example in the AGI definition of hendrycks2025agidefinition, is the arithmetic mean:
@@ 3 A Coherence-Based Measure of General Intelligence
This formulation, denoted as $\mathrm{AGI}_{1}$, assumes that strengths in some faculties can compensate for weaknesses in others, a property known as compensability. Although convenient, such additive aggregation is misaligned with both psychometric evidence and limiting-factor dynamics in complex systems. Severe deficiencies in essential faculties, e.g., reasoning, memory, or perception, often impose system-level bottlenecks that cannot be compensated by strengths elsewhere, causing $\mathrm{AGI}_{1}$ to substantially overestimate holistic capability.
3.1 Generalized Means and Compensability
@@
To relax this compensatory assumption, we generalize Eq. (1) to generalized-mean (power-mean) family (bullen2013handbook), parameterized by a compensability exponent $p\in\mathbb{R}$, where each value of $p$ corresponds to a distinct coherence regime for general intelligence, controlling the degree to which strengths can offset weaknesses:
@@
$\mathrm{AGI}_{p}=\begin{cases}\left(\dfrac{1}{n}\sum_{i=1}^{n}\max\!\left(s_{i},\varepsilon\right)^{p}\right)^{\!\!1/p},&p\neq 0,\\[10.0pt] \left(\prod_{i=1}^{n}\max\!\left(s_{i},\varepsilon\right)\right)^{\!1/n},&p=0,\end{cases}$ (2)
where $\varepsilon>0$ is a small stability constant (set to $10^{-6}$ in all experiments) that prevents numerical collapse when any $s_{i}=0$ and can be viewed as a minimal detectable competence or a noise floor in domain measurement.
The exponent $p$ determines the allowed degree of compensability:
  • $p=1$: arithmetic mean (strongly compensatory),
  • $p=0$: geometric mean (moderately non-compensatory),
  • p=1p=-1: harmonic mean (strongly non-compensatory),
  • $p\to$-$\infty$: minimum operator (strict bottleneck).
@@
As $p$ decreases, $\mathrm{AGI}_{p}$ shifts from measuring average proficiency to measuring coherent sufficiency, requiring balanced competence across all essential dimensions. This reflects the intuition that general intelligence is not characterized by isolated peaks but by the absence of critical weaknesses.
The limiting case $p\to$-$\infty$ yields a fully non-compensatory measure equal to the weakest domain score, but is often too sparse to be practical: the overall score changes only when the minimum dimension improves. Even the harmonic mean (p=1p=-1) can collapse toward zero on uneven profiles. Conversely, permissive regimes such as $p=1$ can mask catastrophic deficits, yielding inflated scores despite near-zero performance in key areas.
These considerations motivate restricting attention to a moderate range, such as $p\in[-1,1]$, which enforces meaningful non-compensatory behavior while preserving sensitivity and avoiding degeneracy. Within this interval, the $\mathrm{AGI}_{p}$ curve provides a concise diagnostic of coherence: models with flatter, higher curves exhibit more uniformly distributed competence and greater robustness to compensability assumptions.
For a single scalar summary of progress toward AGI, one may integrate $\mathrm{AGI}_{p}$ over the chosen $p$-range, yielding a measure that captures both robustness and coherence.
3.2 An Integrated Measure of Coherence: $\mathrm{AGI}_{\mathrm{AUC}}$